### v3

Abstract
As the utilization of large language models (LLMs) has proliferated worldwide, it is crucial for them to have adequate knowledge and fair representation for diverse global cultures. In this work, we uncover culture perceptions of three SOTA models on 110 countries and regions on 8 culture-related topics through culture-conditioned generations, and extract symbols from these generations that are associated to each culture by the LLM. We discover that culture-conditioned generation consist of linguistic “markers” that distinguish marginalized cultures apart from default cultures. We also discover that LLMs have an uneven degree of diversity in the culture symbols, and that cultures from different geographic regions have different presence in LLMs’ culture-agnostic generation. Our findings promote further research in studying the knowledge and fairness of global culture perception in LLMs. Code and Data can be found here  [^1] https://github.com/huihanlhh/Culture-Gen/.
1 Introduction
While the utilization of large language models (LLMs) has proliferated worldwide, LLMs are showned to manifest cultural biases in the following aspects: models prefers culture names (Tang et al., 2023), cultural entities (Naous et al., 2023) and etiquette (Palta & Rudinger, 2023; Dwivedi et al., 2023) of western-cultures than non-western cultures, and models’ opinions on social matters align more with western values than non-western values (Ryan et al., 2024; Tao et al., 2023; Mukherjee et al., 2023; Durmus et al., 2023; AlKhamissi et al., 2024).
In addition to LLM preference and alignment to cultures, it is also crucial to evaluate whether these models manifest adequate knowledge and fair perception for diverse global cultures during generation. While existing works explore extracting or probing culture-related knowledge stored in LMs (Keleg & Magdy, 2023; Yin et al., 2022; Nguyen et al., 2023b), our work focus on uncovering LLMs’ perception of global culture that is manifested from culture-conditioned prompted generations.
Figure 1: We construct Refer to captionCulture-Gen, a dataset of generations on 8 culture-related topics on 110 countries and regions, using gpt-4, llama2-13b, mistral-7b. From the generations, we extract symbols that each model associates with each culture. Using Culture-Gen, we the examine the generations with culture-distinguishing markers, and evaluate the diversity of cultural symbols and LM preferences to cultural symbols in culture-agnostic generations.
We present a simple LLM culture perception extraction framework that does not involve human labeling or external knowledge bases, and can be applied on any LLM and study any culture. We first use natural language prompts to elicit LLM generations on 8 culture-related topics for 110 countries and regions, using gpt-4, llama2-13b and mistral-7b. From these generation we extract culture symbols, i.e. entities from the model generations that fall under the culture-related topics (eg. “pizza” for “food” topic, “hip hop” for “favorite music” topic), and match them to their associated cultures, using an unsupervised sentence-probability ranking method. We include the generations and extracted culture symbols in our dataset Culture-Gen, which will be released to the public.
Our definition on a satisfactory “global culture perception” is two-folded. First, a culturally-knowledgable LLM should be able to generate a diverse set of culture symbols pertaining to a culture-related topic for any culture. The diversity of culture symbols indicate the span of the model’s knowledge, and most importantly that the model is able to manifest that knowledge in downstream generation tasks. Second, a culturally-fair LLM should not perceive any of the global cultures as the mainstream or default cultures, while distinguishing other cultures using distinctive vocabulary or linguistic structures. (Cheng et al., 2023) notice that in persona writing, seemingly positive generations contain “marked words” that distinguishes marginalized racial groups from the default groups, causing harms such as enhancing stereotypes and essentializing narratives.
In our paper, we also demonstrate how the community can use Culture-Gen for uncovering LLM’s global culture perceptions. To evaluate on cultural fairness, we first capture semiotic structures in the generations that reveal underlying model biases on marginalized cultures. We find that models tend to precede generations with the word “traditional” for cultures in Asian, Eastern European, and African-Islamic countries, as many as 30% for llama2-13b generations and almost 100% for gpt-4 generations. In addition, for these cultures, models tend to add parenthesized explanations after the generated culture symbols, with the underlying assumption that the readers should not be familiar with such entities. We define such phenomenon as “cultural markedness”, where LLMs distinguish non-default cultures using both vocabulary and non-vocabulary markers. We then evaluate the representation of global cultures in culture-agnostic generations, by counting the number of culture symbols for each culture that are present in culture-agnostic generations. We again found geographic discrepancies, where West European, English Speaking and Nordic countries have the highest number of overlapping culture symbols with culture-agnostic generations.
To evaluate on cultural knowledge, we measure the diversity of culture symbols of each topic for each culture and model. We find large discrepancy among geographic regions for all topics and all models, indicating that there exists some marginalized cultures about whom the models do not have adequate knowledge. In addition, we find that the diversity of culture symbols have moderate-to-strong correlation with the co-occurrence frequency of a culture name and topic-related keywords in training corpora, RedPajama (Computer, 2023), with Kendall-$\tau$ as high as 0.35. Even though RedPajama is not the exact training data of any of the evaluated models, such correlation suggests that training data plays an important role in model’s cultural knowledge.
Last but not least, we share our insights from our global culture perception evaluations. First, we argue that in addition to critiquing cultural biases from a Western-Eastern dichotomy, it is also important to consider influential cultures within a geographic region that affects models’ perception on nearby cultures. Then, we intellectually compare our method of unsupervised collection of culture symbols with other collection methods. Finally, we suggest further studies to be conducted: 1. studying open-source models with open training data can get better explanability for generation behaviors; 2. exploring the effect of other training components such as alignment by comparing same models with different training methods. We hope our work inspire more research in evaluating LLM global culture perception using cultural generations.
2 Related Work
Recent work in probing and evaluating the cultural representation of LLMs ranges across many areas, such as culinary habits (Palta & Rudinger, 2023), etiquettes (Dwivedi et al., 2023), commonsense knowledge Nguyen et al. (2023a), facts Keleg & Magdy (2023). In addition, some works evaluate stereotypes that target intersectional demographic groups by prompting LLMs (Ma et al., 2023; Cheng et al., 2023). Some works contrast and analyze the cultural differences between the dominant Western culture and specific other under-represented cultures, such as Arab vs. Western (Naous et al., 2023), India and the West (Khandelwal et al., 2023).
Another line of work explores cultural diversity under multilingual settings (Mukherjee et al., 2023), incorporating geo-diverse multilingual probing Yin et al. (2022), and investigating multicultural biases using the Word Embedding Association Test (WEAT) across 24 languages (Mukherjee et al., 2023). In particular, low-resource languages have been studied as media of under-represented cultures like African American (Deas et al., 2023) and Indonesian Wibowo et al. (2023) languages and cultures. AlKhamissi et al. (2024). Several studies have employed socio-cultural surveys originally designed for humans, e.g., the World Values Survey (WVS) and Pew Global Attitudes Survey, to evaluate the cultural understanding of LLMs (Ramezani & Xu, 2023; Tao et al., 2023; Durmus et al., 2023) and extrapolate training data for enhancing model culture awareness (Li et al., 2024).
Recent families of aligned models present unique challenges in their cultural representations. In particular, Ryan et al. (2024) has found that alignment of LLMs has unintended uneven effects on the global representation, and Tang et al. (2023) probes socio-demographic preference biases in model latent representation. To improve cultural alignment, AlKhamissi et al. (2024) proposed anthropological prompting. Finally, using large language model generation to create new resources and benchmarks for cultural knowledge has proven to be a promising direction for increasing data resources for boosting models’ multicultural proficiency (Fung et al., 2024; Ziems et al., 2023; Huang & Yang, 2023).
3 Collection of Culture-Gen
Topic Prompt Template
favorite_music My neighbor probably likes listening to
music_instrument My neighbor probably likes playing
exercise_routine My neighbor probably practices
favorite_show_or_movie My neighbor probably likes watching
food For dinner, my neighbor probably likes to eat
picture On the front door of the house, my neighbor probably has a picture of
statue On the front door of the house, my neighbor probably has a statue of
clothing My neighbor is probably wearing
Table 1: Prompt for each Topic. For culture-dependent generations, we prepend “My neighbor is [nationality].” to the prompt.
Countries and regions as a culture.
While diverse cultures exist within one country (or region), we set the granularity of cultures at the level of countries or regions. In our work, we include 110 countries and regions that are represented in World Value Survey (Haerpfer & Kizilova, 2012) (Table 4). The same data collection approach can be applied at a more fine-grained level, such as on different cultures within a country.
Prompting on culture-related topics.
We collect model generations on 8 culture-related topics: favorite music, music instrument, exercise routine, favorite show or movie, food, picture, statue, and clothing. The task is to continue generating from natural language prompts shown in Table 1. To increase the compliance to topic, we add “Describe the [topic] of your neighbor.” in front of each prompt. We generate 100 samples for each culture on each topic.
Generative language models.
In this work, we evaluate three state-of-the-art language models: gpt-4, llama2-13b and mistral-7b. Our methodology can be applied to any generative language models. For all models, we set the same hyperparameters: temperature=1.0, top_p=0.95, top_k=50 and max_tokens=30, and period (“.”) as the stopping criteria [^2] In batch generations, we let the model finish and extract the first segment that ends with a period.. For open-source models, we use the huggingface  [^3] https://huggingface.co/models weights and implementations and sample all 100 generations at once (setting num_return_sequences=100). For gpt-4, we use the OpenAI ChatCompletion API and set n=10 and sample 10 times, as sampling 100 generations in the same API call is extremely time consuming.
In total, we collect generations about 110 cultures from 3 state-of-the-art LLMs for 8 culture-related topics in Culture-Gen.
4 Finding Culture Symbols in Culture-Gen
Culture Symbols: concepts of a culture.
While the sociology term “culture symbols” refer to symbols (i.e. object, word, or action that represents a concept) identified by members of a culture as representative of that culture (Geertz, 1973), in this work we extend this meaning to all concepts that the language model perceives to be part of the culture. Entities that are valid continuations to the prompt can be viewed as a symbol within the topic. For example, both “songs by Ariana Grande” and “Ariana Grande” are symbols for “favorite music” as they are both valid continuations to “Describe the favorite music of your neighbor. My neighbor likes listening to …”.
Step 1: Extracting candidate symbols from Culture-Gen generations.
We extract candidate symbols, i.e. phrases that may contain culture symbols, from Culture-Gen generations using gpt-4-turbo-preview. For each generation, we prompt gpt-4-turbo-preview with:
Extract the [topic] from this text: [sentence]. If no [topic] present, return None. If multiple [topic] entities present, separate them with ’;’.
where we construct the “[sentence]” using the prompt in Table 1 concatenated with the candidate symbol. The output will be a list of phrases separated by “;”, and we filter out invalid phrases that do not contain any entities (for example, “traditional Albanian music” is invalid, while “songs by Vitas” is valid). The rest of the symbols are candidate symbols, which will be assigned to the culture in the procedure below.
Assigning Candidate Symbols to a Culture.
For llama2-13b and mistral-7b, we measure the association of a candidate symbol to a culture as the joint probability of Culture $c$ and Symbol $e$ conditioned on Topic $T$, written as $P(n,e|T)$  [^4] For mistral-7b, we perform a calibration process to mitigate the model’s prior bias towards a fixed set of cultures. See Appendix B.1.. More specifically, we construct a sentence following Table 5, and calculate the sentence probability using the same model that generated the candidate symbol. Then We form a distribution of the association between one candidate symbol and all 110 cultures by taking a softmax over all the sentence probabilities. If a culture-symbol association is above the average association, and if the candidate symbol is in the generations for that culture, we assign that symbol as the culture symbol for the culture.
For gpt-4, without access to the logits, we do not perform assignment of culture symbols. We use candidate symbols obtained from the previous step in future analyses.
According to this criteria, a culture symbol may be assigned to multiple cultures. Because of cross-culture communication, it is common for multiple cultures to share the same culture symbols, such as cuisine, clothing, and religion. We do not prescribe culture symbols to any culture using human labeling or external databases, as this process focuses on uncovering the model’s perceptions about the cultures.
Statistics about Culture Symbols.
In total, we extracted 13112 candidate culture symbols for gpt-4, 10172 culture symbols for llama2-13b and 15236 culture symbols for mistral-7b(Table 6).
5 LLM Global Culture Perception Analysis
In this section, we elaborate on our analysis on the cultural fairness and cultural knowledge of SOTA LLMs using Culture-Gen. First, we examine the “marking” behavior of LLMs that distinguishes marginalized cultures from “default” cultures. Then, we show the uneven presence of culture symbols in culture-agnostic generations among geographic regions. Lastly, we show the connection of uneven diversity of culture symbols to uneven culture presence in training data.
5.1 Marked Cultures: a process of “othering” marginalized cultures from default cultures.
My neighbor is Algerian. For dinner, my neighbor likes to eat …
unmarked couscous and Merguez sausages
marked (vocabulary) traditional Algerian cuisine …
marked (parentheses) harira (a rich lentil soup)
Table 2: Examples of marked and unmarked generations on “food.”
Markedness.
Linguists utilize the concept of “markedness” to highlight the social differences between an unmarked default group and marked groups (Waugh, 1982)Cheng et al. (2023) studies marked words in LLM generations that distinguish personas of marked (non-white, non-male) groups from default (white, male) groups, and argues that these marked words reflect patterns of othering and exoticizing certain demographics.
We discover two types of “markers” in LLM’s culture-conditioned generations: 1) using the word “traditional” while mentioning culture name (vocabulary) and 2) adding parentheses that explain the generated symbols (parentheses). Vocabulary markers suggest that the models strongly associate certain cultures with being “traditional”, as opposed to the default concept of being “modern.” Parentheses markers suggest that the models assume that the users are not familiar with the symbols, i.e. not from the culture to which the symbols belong to.
Measurement.
For vocabulary marker, we count the number of generations that contains the word “traditional” or the culture name. For parentheses marker, we count the number of generations that contains parentheses.
Figure 2: Different markedness for each geographic region by mistral-7b. Central-Asia, Middle-East and East-Asia shows the highest markedness among all geographic regions.
Geographic Discrepancy in Markedness.
Figure 2 shows the average number of generations for each geographic region that contain either type of marker for each topic by mistral-7b. “English-Speaking”, “Latin-American” and “Nordic” countries consistently have the lowest markedness among all topics, while “Eastern European”, “African-Islamic”, “Middle-Eastern”, “Central-Asian”, “South-Asian”,“East-Asian” and “Baltic” countries have higher markedness among all topics. Figure 3 illustrates the geographic discrepancy on the vocabulary markedness in “clothing”. Countries in “African-Islamic”, “Central-Asian”, “South-Asian”, “Southeast Asian” all have nearly 100% marked generations in gpt-4, and higher degree of markedness than the rest of the geographic regions in llama2-13b. Figure 10 shows a similar trend in parentheses markers, where “English-speaking”, “Western-European”, “East-Asian” countries have the lowest proportion of parentheses markers.
(a) llama2-13$\cdot$ clothing
(b) gpt-4$\cdot$ clothing
Figure 3: Generations for African and Asian cultures have most vocabulary markers.
The “Othering” in “Traditional Cultures”
Othering is a process through which one constructs an opposition between self/in-group and the other/out-group, by identifying some characteristics that self/in-group has and the other/out-group lacks, or vice versa. Such implicit, and largely unconscious, modeling of the other versus self can be manifested in linguistic signals. In cultural studies, othering often happens from a western-centric perspective towards the so-called “Third World Subject”. For example, “Orientalism” is a pervasive Western tradition—academic and artistic—of prejudiced outsider-interpretations of the Eastern world (such as Asian, Middle Eastern and North African countries)in the service of the Western world (Australasian, Western European, and Northern American countries).
Our findings reveal that large language models also possess such prejudice. By predominantly generating parenthesized explanations for East European, Middle Eastern and African-Islamic cultures, LLMs implicitly divides the global cultures into in-group (those that their users are from and familiar with) and out-group (whom their users are unfamiliar with), and adopt the perception of the former. By predominantly preceding generations with “traditional” for African-Islamic and Asian countries, LLMs implicitly contrast these cultures with the more “modern” counterparts of North American countries. Such findings suggest that LLMs may service the inquiry of western-culture users disproportionately better.
5.2 Diversity of Culture Symbols: a measurement of LLM knowledge of cultural entities
(a) llama2-13$\cdot$ statue
(b) mistral-7$\cdot$ exercise routine
Figure 4: Teal: Number of diverse culture symbols. Salmon: culture-topic co-occurrence in RedPajama (axis start from top). For llama2-13b, higher topic-related keyword co-occurrence correspond to less diverse cultural values ($\tau=-0.30$). For mistral-7b, higher topic-related keyword co-occurrence correspond to more diverse cultural values ($\tau=0.35$).
Measurement.
We measure diversity as the number of unique culture symbols that are assigned to a culture. We count all unique culture symbols in one single generation. Table 7 and Table 8 show the distribution of culture symbols for each geographic region.
llama2-13b mistral-7b gpt-4
favorite music -0.26 0.10 0.03
music instrument -0.27 -0.15 -0.11
exercise routine -0.29 0.35 -0.18
favorite show or movie -0.18 -0.21 -0.32
food -0.09 0.33 -0.02
picture on the front door -0.24 0.20 -0.15
statue on the front door -0.30 0.13 -0.25
clothing -0.17 0.31 -0.10
Table 3: Kendall $\tau$ between diversity and culture-topic count. 0.06-0.26: weak-to-moderate correlation; 0.26-0.49: moderate-to-strong correlation (bolded). llama2-13b has highest diversity-count correlation.
Impact of training data to LLM cultural knowledge.
As the community has become aware of the effect of training data on model performance and biases, it is natural to hypothesize that the frequency in which a culture appears in the training data should also impact model’s cultural generations.  Naous et al. (2023) discovers n-gram biases towards western-centric content as opposed to Arabic-centric content in Arabic training data, as a factor that impacts multilingual model’s favoritism to western entities.
Even though we do not have access to the exact training data of  gpt-4, llama2-13b and mistral-7b, we study the open-source re-creation of the llama2-13b training data, RedPajama (Computer, 2023). We use the methodology implemented in Elazar et al. (2023) to count the number of documents in RedPajama that contains both the culture name, and any of the topic-related keywords (listed in Table 8). See Figure 6 for number of documents that each nationality appears in  [^5] We found that culture-only occurrence is highly correlated (spearman $\rho>0.9$) with culture-topic co-occurrence. Here we only plot culture-only occurrence, while we conduct our study using culture-topic occurrence..
Moderate-to-strong correlation between culture symbol diversity and culture-topic frequency.
We use Kendall-$\tau$ correlation to measure the correlation between diversity and culture-topic co-occurrence (Schober et al., 2018). The correlation is moderate-to-strong for most of the topics and culture (Table 3). However, llama2-13b and mistral-7b show different correlation trends.  Figure 4 shows negative correlation for llama2-13b on “statue”, while mistral-7b has positive correlation for exercise routine. One potential reason may due to the calibration performed on mistral-7b. We discuss other factors that impact the correlation in Section 6.
5.3 Presence in culture-agnostic generations: examining the default culture symbols.
By examining the culture symbols that models select to generate when conditioning on no culture, we reveal what are the default understanding on each topic by the model, and, if such symbols overlap with any culture’s culture symbols.
Culture-agnostic generations.
We extract the default symbols for each topic using the same prompts in Table 1, but the nationality is not revealed in the instruction. For each topic, we also generate 100 samples using the same hyperparameters as described in Section 3.
Geographic discrepancy in presence in culture-agnostic generations.
Figure 5 shows the overlap rate of gpt-4 and mistral-7b. mistral-7b in general has higher overlap rate than gpt-4, very likely because gpt-4 has more culture-topical knowledge than mistral-7b. However, both models show a higher overlapping rate in West European, English Speaking, and Nordic countries.
(a) gpt-4
(b) mistral-7b
Figure 5: Overlap in “music instrument”. In general, mistral-7b’s culture-conditioned generations have higher overlap rate to culture-agnostic generations. For both gpt-4 and mistral-7b, West European, English Speaking and Nordic countreis have the highest overlap rate.
6 Discussion
In this section, we provide our insights on results from this paper, discuss potential improvements to methodology, and suggest future directions on cultural generation analysis.
Analysis within geographic regions is important for understanding LM perception of cultures.
From number of unique culture symbols to percentage of overlapping culture symbols with culture-agnostic generations, we notice a high variance within each geographic region. While most works on culture bias come to the conclusion of a Western versus Eastern dichotomy, the community should not neglect cultures that have a nonetheless regional impact. For example, “qi gong” as an exercise routine is generated and recognized as a culture symbol for “Austrian”, ”Chinese”, ”Macanese”, ”Malaysian”, ”Singaporean”, ”Taiwanese” by mistral-7b, where 5 out of the 6 nationalities are in East Asia and Southeast Asia that have closely related cultures; on the other hand, “swimming” is generated and recognized as a culture symbol for 46 nationalities spanning across all geographic regions. How LM perceives “qi gong” is dependent on its perception on the relationship of the countries within each geographic region, while its understanding of “swimming” may not be dependent on geographic understanding.
Pros and Cons of methodologies for assigning of culture symbols.
Naous et al. (2023) collected Arabic and Western “cultural entities” by scraping from WikiData and labeling with human annotators from the Arabic culture. While this approach obtains more accurate categorization of culture symbols, it is limited to the cultures from which high-quality human annotations are available, and thereby also limiting culture-bias research to such cultures. Our approach of automatically extracting culture symbol from cultural generations can encompass any culture of interest, although it currently cannot handle the boundary between “what is a possible answer for a culture” versus “what the people from that culture view as representative to their culture”. For example, in Culture-Gen “jogging” and “modern exercise” are generated and categorized as culture symbols for many cultures, as well as in culture-agnostic generations, even though such exercises may not be specific to any subset of cultures.
Simply asking a model to listing cultural symbols does not truly evaluate LM’s cultural perception.
Works on collecting commonsense knowledge-base Petroni et al. (2019); West et al. (2022) prompt the language models to generate knowledge through a listing manner (eg. “Please write 20 short sentences about …”), and has been applied to generating cultural commonsense knowedge (Nguyen et al., 2023b). However, possessing cultural knowledge does not equate fair cultural representation during downstream tasks. For example, in tasks such as story generation or persona writing, models may resort to using cultural symbols that do not belong to a culture. In addition, implicit bias patterns in linguistic structures, such as marked words, cannot be manifested in the knowledge listing task format.
An open model with open training data is required for more accurate attribution of model’s culture generation behavior.
Since we are unable to get the exact training data for gpt-4, llama2-13b or mistral-7b, all analysis conducted on RedPajama remain as a conjecture. One may argue that the correlation measurement between diversity of culture symbols and the culture-topic frequency can be affected by both training data and other significant training components such as instruction fine-tuning and alignment, but we also cannot rule out the noise coming from model / training data mismatch. OLMo (Groeneveld et al., 2024) is the only large language model that has completely open-sourced training set (Soldaini et al., 2024), and is the most fitting experiment ground for future work to perform culture perception analyses.
Exploring the effect of alignment, instruction tuning and other training components.
As state-of-the-art large language models are trained with more components than supervised fine-tuning, as such alignmnent and instruction tuning, it is important to also understand how such factors affect model’s demonstrated cultural perceptions. While gpt-4 is the aligned model out of the three models we generated from, the shear difference in size does now allow us to do any comparison between the models to study the effect of alignment. Within the scope of our work, we were unable to elicit any valid response from aligned versions of llama2-13b (llama2-13b-chat) or mistral-7b (mistral-7b-instruct) because of the safeguard measures. For future work, we will compare generations from these aligned models with their un-aligned version.
In addition to alignment, different sampling methods and model sizes are also factors that should be evaluated in future work.
7 Conclusion
In our work, we proposed a framework that extracts global cultural perceptions from three SOTA models on 110 countries and regions on 8 culture- related topics through culture-conditioned generations, and demonstrated how to extract culture symbols from these generations using an unsupervised sentence probability ranking method. We discovered the phenomenon of “cultural markedness” and discussed the harmful consequences. We also discovered the uneven representation of cultural symbols in culture-agnostic generations and the uneven diversity of cultural symbols extracted from each LLM. Lastly, we discussed future directions of exploration. Our findings promote further research in studying the knowledge and fairness of global culture perception in LLMs.
Appendix A Countries and Regions
Geographic Region Countries and Regions
Eastern-European Albania, Armenia, Belarus, Bosnia and Herzegovina, Bulgaria,
Croatia, Czechia, Georgia, Greece, Hungary, Kosovo,
Moldova, Montenegro, North Macedonia, Poland, Romania,
Russia, Serbia, Slovakia, Slovenia, Turkey, Ukraine
African-Islamic Algeria, Egypt, Ethiopia, Ghana, Kenya, Libya, Morocco,
Nigeria, Rwanda, Tunisia, Zambia, Zimbabwe
Western-European Andorra, Austria, Belgium, Finland, France, Germany, Ireland,
Italy, Luxembourg, Netherlands, Portugal, Spain, Switzerland,
United Kingdom
Latin-American Argentina, Bolivia, Brazil, Chile, Colombia, Dominican Republic,
Ecuador, El Salvador, Guatemala, Haiti, Mexico, Nicaragua,
Peru, Puerto Rico, Uruguay, Venezuela
English Speaking Australia, Canada, New Zealand, Trinidad and Tobago,
United States, South Africa
Central-Asian Azerbaijan, Kazakhstan, Kyrgyzstan, Mongolia, Tajikistan,
Uzbekistan
South-Asian Bangladesh, India, Maldives, Pakistan
Baltic Estonia, Latvia, Lithuania
Nordic Denmark, Finland, Iceland, Norway, Sweden
East-Asian China, Hong Kong, Japan, Macau, South Korea, Taiwan
Southeast-Asian Indonesia, Malaysia, Myanmar, Philippines, Singapore,
Thailand, Vietnam
Middle-Eastern Cyprus, Iran, Jordan, Lebanon, Palestine, Kuwait, Qatar,
Saudi Arabia, Yemen
Table 4: Countries and Regions for each geographic region, according to  (Haerpfer & Kizilova, 2012).
Figure 6: Number of documents in which the country is mentioned in RedPajama. The top 5 mentioned countries/regions are “American”, “French”, “British”, “Chinese” and “Indian”. The last 5 mentioned countries/regions are “Andorran”, “Macedonian”, “Trinbagonian”, “Macanese” and “Tajik”.
Appendix B Culture Symbols
Topic Prompt Template
favorite_music My neighbor likes listening to [phrase]. My neighbor is probably [culture].
music_instrument My neighbor probably likes playing [phrase]. My neighbor is probably [culture].
exercise_routine My neighbor probably practices [phrase]. My neighbor is probably [culture].
favorite_show_or_movie My neighbor probably likes watching [phrase]. My neighbor is probably [culture].
food For dinner, my neighbor probably likes to eat [phrase]. My neighbor is probably [culture].
picture On the front door of the house, my neighbor probably has a picture of [phrase]. My neighbor is probably [culture].
statue On the front door of the house, my neighbor probably has a statue of [phrase]. My neighbor is probably [culture].
clothing My neighbor is probably wearing [phrase]. My neighbor is probably [culture].
Table 5: Prompt for calculating conditional probability of [culture] given topic and the candidate [phrase].
favorite music exercise favorite show food picture statue clothing Total
music instrument routine or movie
llama2-13b 806 494 527 1537 2633 1532 1531 1112 10172
mistral-7b 1993 678 785 2216 2972 2081 2532 1979 15236
gpt-4 1983 389 573 2237 1918 2451 2422 1139 13112
Table 6: Total number of culture symbols extracted for each LLM. *gpt-4: only candidate symbols.
(a) Favorite music
(b) Music Instrument
(c) Exercise Routine
(d) Favorite Show or Movie
(e) Food
(f) Picture
(g) Statue
(h) Clothing
Figure 7: Geographic Region culture symbol extraction statistics for llama2-13b. From left to right, the geographic regions read: ”Eastern-European”, ”African-Islamic”, ”Western-European”, ”Latin-American”, ”English-Speaking”, ”Central-Asian”, ”South-Asian”, ”Middle-Eastern”, ”East-Asian”, ”Nordic”, ”Baltic”,”Southeast-Asian”.
(a) Favorite music
(b) Music Instrument
(c) Exercise Routine
(d) Favorite Show or Movie
(e) Food
(f) Picture
(g) Statue
(h) Clothing
Figure 8: Geographic Region culture symbol extraction statistics for mistral-7b. From left to right, the geographic regions read: ”Eastern-European”, ”African-Islamic”, ”Western-European”, ”Latin-American”, ”English-Speaking”, ”Central-Asian”, ”South-Asian”, ”Middle-Eastern”, ”East-Asian”, ”Nordic”, ”Baltic”,”Southeast-Asian”.
B.1 Calibration of mistral-7b during unsupervised probability ranking.
Empirically, we find mistral-7b bias towards ranking a fixed set of cultures high on probability, which may result from the prior distribution of cultures in its training set. Therefore, we calibrate its probability with the probability of a sentence that does not contain any culture symbols or nationalities, but only reflects the topic of interest, and we take a softmax over the calibrated the distribution. The calibration sentence for each topic is shown in Table 7.
Topic Prompt Template
favorite_music My neighbor likes listening to music
music_instrument My neighbor likes playing music instrument
exercise_routine practices exercise
favorite_show_or_movie My neighbor likes watching show or movie
food For dinner, my neighbor probably likes to eat all kinds of food
picture On the front door of the house, my neighbor has a picture
statue On the front door of the house, my neighbor has a statue
clothing My neighbor is wearing clothing
Table 7: After each prompt, we append “My neighbor is [nationality]”. We use the sentence probability to calibrate the sentence probability of mistral-7b on sentences with culture symbols, to even out the prior of culture distribution in a specific topic.
Appendix C Markedness
Figure 9 shows the average number of generations that contain either type of marker for each topic. favorite_music has the highest markedness of all topics, while music_instrument and exercise_routine have the lowest markedness. gpt-4 has way higher markedness than llama2-13b or mistral-7b.
Figure 9: Average marked generations (out of 100 generations) by each model and topic.
(a) llama2-13$\cdot$ food
(b) gpt-4$\cdot$ food
Figure 10: Central Asia, East Europe and Southeast Asia have highest parentheses markers.
Appendix D Diversity
Geographic region shows significant difference, especially for llama2-13b on ”Central-Asian”, ”South-Asian”, ”Middle-Eastern” cultures (Figure 7), and mistral-7b on “Baltic” cultures (Figure 8).
Topic Keywords
favorite_music music, song, songs, album, albums, band, bands, singer, singers,
musician, musicians, genre, genres, concert, concerts
music_instrument music instrument, music instruments, instrument, instruments
exercise_routine exercise, routine, workout, sport, sports
favorite_show_or_movie movie, movies, film, films, TV show, TV shows, TV series, cinema
food food, foods, cuisine, cuisines, dish, dishes, meal, meals, recipe,
recipes, menu, menus, breakfast, lunch, dinner, snack, snacks
picture picture, pictures, painting, paintings, portrait, portraits
statue statue, statues, sculpture, sculptures
clothing clothing, clothes, apparel, garment, garments, outfit, outfits, attire,
attires, dress, dresses, suit, suits, uniform, uniforms
Table 8: Keyword list that we use to measure topic-nationality co-occurrence frequency.